# Bound excited states of Fröhlich polarons in one dimension

### v1

@@ Abstract
The one-dimensional Fröhlich model describing the motion of a single electron interacting with optical phonons is a paradigmatic model of quantum many-body physics. We predict the existence of an arbitrarily large number of bound excited states in the strong coupling limit and calculate their excitation energies. Numerical simulations of a discretized model demonstrate the complete amelioration of the projector Monte Carlo sign problem by walker annihilation in an infinite Hilbert space. They reveal the threshold for the occurrence of the first bound excited states at a value of $\alpha\approx 1.73$ for the dimensionless coupling constant. This puts the threshold into the regime of intermediate interaction strength. We calculate the spectral weight and the number density of the dressing cloud in the discretized model.
Studies of the polaron, describing the motion of an electron in a polarizable medium, go back almost a century to the works of Landau [1] and Pekar [2]. Formalised by Fröhlich [3], the polaron model is one of the most studied problems of quantum many-body theory [4] with real-world relevance for electronic material properties [5]. Calculations of the ground state energy, the dispersion relation and also the structure of the excitation continuum have been achieved at great accuracy and become an important test bed for analytical [6] and numerical [7, 8, 9, 10] techniques. Thus it may seem surprising that evidence for bound excited states below the excitation continuum has emerged only very recently [11].
The proof of the existence of arbitrarily many bound excited states in the strong-coupling limit of the three-dimensional model [11] demonstrates that previous approaches have missed key features of the Fröhlich polaron at intermediate and strong coupling. Previous numerical studies have failed to produce evidence for bound excited states in the three-dimensional case [8]. Studies of one-dimensional Fröhlich polaron models found excitations of the electron as resonances at high energies within the phonon continuum [12, 13]. In contrast, for the Holstein model of a small polaron, the existence of a bound excited state below the continuum is well known [14, 15, 16].
In this Letter we calculate the energies of the bound excited states of the phonon dressing cloud for the Fröhlich model in one dimension in the strong coupling limit, and provide numerical evidence for them in the weak to intermediate coupling regime applying Full Configuration Interaction Quantum Monte Carlo (FCIQMC) [17, 18, 19]. The first bound excited state appears at $\alpha\approx 1.73$, which is in a crossover regime between weak and intermediate coupling strength with a spectral weight of $\approx 0.2$, suggesting that these states may be accessible in experiments.
@@
The model – The one-dimensional Fröhlich model describes an electron interacting with lattice vibrations in one dimension. The original three-dimensional Fröhlich Hamiltonian [3] was generalized to arbitrary dimensions in Refs. [32, 33]. The Hamiltonian in the one-dimensional case is
@@
$-\hbar\omega_{\mathrm{LO}}\sqrt{\frac{2\alpha}{L/l_{0}}}\sum_{k}\left(\hat{a}_{k}e^{ik\hat{r}}+\hat{a}^{\dagger}_{k}e^{-ik\hat{r}}\right),$ (1)
where $\hat{p}$ and $\hat{r}$ are the electron momentum and position, $\hat{a}^{\dagger}_{k}$ and $\hat{a}_{k}$ are the phonon creation and annihilation operators, $\omega_{\mathrm{LO}}$ is the frequency of longitudinal optical phonons (taken to be constant), $m$ is the band mass of the electron, and $\alpha$ is a dimensionless coupling constant. $L$ is the length of the system with periodic boundaries (eventually taken to be infinite), and $l_{0}=\sqrt{{\hbar}/{2m\omega_{\mathrm{LO}}}}$ is a natural unit of length.
@@
Excitation energies in the strong coupling limit – For large $\alpha$, the discrete excited eigenvalues of $\hat{H}$ in Eq. (Bound excited states of Fröhlich polarons in one dimension) at total momentum $P=0$ can be obtained analytically, and are given by
@@
As is well-known [37, 38, 39], the first term, $-\alpha^{2}/3$, represents the semiclassical energy obtained by minimizing the corresponding Pekar functional with respect to the electron wave function $\psi(x)$ and the classical polarization field $\phi(x)$. The Pekar functional (given in dimensionless units integrating over $x=r/l_{0}$)
@@
is obtained as an expectation value of the Hamiltonian of Eq. (Bound excited states of Fröhlich polarons in one dimension) with respect to a product of the electron wave function $\psi(x)$ and a coherent state for the phonon field with amplitude $\varphi(x)$. The physical picture is that, at large coupling, the polarization field effectively becomes classical, and the electron occupies the ground state in the corresponding effective potential. The Pekar functional has the explicit minimizer $\psi^{\rm P}(x)=\sqrt{\alpha/2}\,\mathrm{sech}(\alpha x)$ for the electron wave function and $\varphi^{\rm P}(x)=\sqrt{2\alpha}\psi^{\rm P}(x)^{2}$ for the polarization field. The other terms in Eq. (2) involving the ($\alpha$ independent) frequencies $\omega_{j}$ quantify the quantum fluctuations of the field around its classical value [40, 41, 42, 43, 44, 11]. Letting the electron take the instantaneous ground state in a varying field, they are obtained by expanding $\min_{\psi}\mathcal{E}^{\rm P}[\psi,\varphi^{\rm P}+{\varphi}]$ to second order in ${\varphi}(x)$, with the result
@@
$-\frac{\alpha^{2}}{3}+\|\text{Im}({\varphi})\|^{2}+\langle\text{Re}({\varphi}),\hat{\omega}^{2}\text{Re}({\varphi})\rangle.$ (4)
Here the Hessian operator takes the simple form
$\hat{\omega}^{2}=1+8\alpha\partial\psi^{\rm P}(-\partial^{2}+\alpha^{2})^{-2}\psi^{\rm P}\partial,$ (5)
@@
where $\psi^{\rm P}(x)$ is understood as a multiplication operator, and $\partial=d/dx$. By scaling, the eigenvalues of $\hat{\omega}$ are independent of $\alpha$. We denote them by $10=\omega_{0}<\omega_{1}\leq\dots<$ 1. There is a single zero mode, $\omega_{0}=0$, resulting from translation invariance. The lowest non-vanishing eigenvalues are given by $\omega_{1}\approx 0.647$ and $\omega_{2}=\sqrt{7/9}\approx 0.882$. For details, we refer to Appendix A in the End Matter.
@@
This leads to a contribution to the ground state energy of the form $\frac{1}{2}\sum_{j\geq 0}\omega_{j}=\infty$ from the zero-point motion. To obtain the finite quantum corrections, we need to subtract the (infinite) energy of the fluctuations in the absence of the electron, however, leading to the corrected expression $\tfrac{1}{2}\sum_{j\geq 0}(\omega_{j}-1)$, hence justifying Eq. (2) for $j=0$. Numerically, one finds $\tfrac{1}{2}\sum_{j\geq 0}(\omega_{j}-1)\approx-0.95$. For $j\geq 1$, the excitation energies correspond to excitations of different oscillators in Eq. (6). Since $\omega_{j}>\tfrac{1}{2}$ for $j\geq 1$, only single excitations can appear below the continuum. The zero mode $j=0$ needs to be discarded for the calculation of the excitation spectrum of Eq. (Bound excited states of Fröhlich polarons in one dimension) due to the restriction to zero total momentum.
Lee-Low-Pines transformation – Using a unitary transformation due to Lee, Low, and Pines amounting to a frame transformation [45], the electron coordinate dependence of the Fröhlich polaron Hamiltonian of Eq. (Bound excited states of Fröhlich polarons in one dimension) can be eliminated and the Hamiltonian rewritten in terms of the conserved total momentum $P$:
@@
$-\hbar\omega_{\mathrm{LO}}\sqrt{\frac{2\alpha}{L/l_{0}}}\sum_{k}\left(\hat{a}^{\dagger}_{k}+\hat{a}_{k}\right).$ (7)
This equivalent Hamiltonian acts only on a Fock space of phonons for a given value of the total momentum $P$ and is more convenient for numerical calculations. We take the system size $L$ finite, which discretizes the phonon modes with the values of $k$ restricted to multiples of $2\pi/L$. A finite number of modes $M$ used in numerical calculations results in a momentum cutoff of $k_{c}=\pi(M-1)/L$. Even so, the phonon Hilbert space dimension is still formally (and practically) infinite, as an arbitrary number of phonons can reside in each mode. Technically, the number of phonons in each mode is limited to be $<256$ by the 8-bit number type used for representing the Fock basis. We perform calculations with up to $M=73$ and Hilbert space dimension $256^{M}\approx 6\times 10^{175}$. In the following, we will restrict the analysis to $P=0$.
FCIQMC – The algorithm samples the ground state of $\hat{H}_{\mathrm{LLP}}$ by repeatedly applying the operator $1-\text{d}\tau(\hat{H}_{\mathrm{LLP}}-S)$ to a coefficient vector. The shift parameter $S$ is updated to keep the norm (the number of ‘walkers’) approximately constant [21], and therefore provides an estimator of the eigenstate energy after equilibration. The operation is performed exactly for coefficients with a modulus above a given threshold and stochastically for the rest. Combined with stochastic vector compression [46], this implements a semistochastic algorithm, where part of the Hamiltonian matrix is treated deterministically while the rest is sampled stochastically [18, 25]. Crucially, contributions of opposite sign to a coefficient can annihilate and thus reduce sign noise. At sufficiently large walker number, the sign of the dominant coefficients becomes stable by a spontaneous symmetry-breaking process [28, 47] and the sign problem is mitigated. The residual noise leads to stochastic errors and also causes a small population control bias [48, 49, 22], which we control by increasing the walker number until undetectable levels of the bias are seen that are smaller than the stochastic error bars.
If both the box size $L$ and the momentum cutoff $k_{c}$ are large enough, the results obtained using the discretized model approach the continuum limit. We find that the energies converge exponentially in $L$, but only algebraically in $k_{c}$ (see Supplemental Material [50]).
Ground state energy – In the weak coupling limit, the ground state energy has been found analytically using perturbation theory [51, 33] and the path integral formalism [52]. For $P=0$, the ground state energy is
@@
Figure 1 shows the transition from the weak coupling regime, through an intermediate coupling regime around $\alpha\approx 2$ where both expansions approximately overlap, into the strong coupling regime as the coupling strength $\alpha$ increases. While there is no phase transition separating the weak and strong coupling regimes, different physical pictures apply [53]. Our FCIQMC approach allows us to probe both regimes and the transition region comfortably. At $\alpha=2$ we estimate the discretization error to be around 1% and about an order of magnitude larger than the stochastic error [50].
Excited state energies – In order to calculate excited states numerically with FCIQMC, the computation is initialized with multiple starting vectors, which are then orthogonalized at every step (or a specified number of steps) of the computation using the Gram-Schmidt procedure. This ensures that the $n^{\text{th}}$ vector converges to the $n^{\text{th}}$ eigenvector, and the $n^{\text{th}}$ shift acts as an estimator for the $n^{\text{th}}$-lowest eigenvalue [19].
Even though the ground state calculation with the Hamiltonian (Bound excited states of Fröhlich polarons in one dimension) is free of the Monte Carlo sign problem, excited state calculations introduce the sign problem [28, 29] due to the necessary existence of nodes in the excited state wave functions [19]. The sign problem manifests itself at low walker number in fluctuating signs of all state vector coefficients and an average value of the shift $\langle S\rangle$ that is lower than the correct energy [17, 31]. Increasing the walker number above a critical value, the shift average stabilizes and becomes independent of the walker number, indicating that the sign problem is overcome. For the studied range of $\alpha$ values we find that $10^{7}$ walkers is sufficient, indicating that the sign problem is relatively weak in this case [31]. This is remarkable as the unconstrained phonon occupation makes the Hilbert space of the Fröhlich Hamiltonian (Bound excited states of Fröhlich polarons in one dimension) infinite. For details, see Appendix B in the End Matter.
Figure 2: The excitation energies of the first four excited states, with zero total momentum, calculated using FCIQMC for various coupling strengths and box sizes (error bars and symbols). The first bound excited state appears at $\alpha\approx 1.73$, in the intermediate coupling regime. The dashed horizontal line corresponds to the analytical result for the first discrete excitation energy at large $\alpha$. The shaded region shows the phonon continuum for $L\to\infty$. All states other than those labeled as odd are even under a parity transformation. Energies for $L=8l_{0}$ are only shown in the inset.

### v2

@@ Abstract
The one-dimensional Fröhlich model describing the motion of a single electron interacting with optical phonons is a paradigmatic model of quantum many-body physics. We predict the existence of an arbitrarily large number of bound excited states in the strong coupling limit and calculate their excitation energies. Numerical simulations of a discretized model demonstrate the complete amelioration of the projector Monte Carlo sign problem by walker annihilation in an infinite Hilbert space. They reveal the threshold for the occurrence of the first bound excited states at a value of $\alpha\approx 1.73$ for the dimensionless coupling constant. This puts the threshold into the regime of intermediate interaction strength. We find a significant spectral weight and increased phonon number of the bound excited state at threshold.
I Introduction
Studies of the polaron, describing the motion of an electron in a polarizable medium, go back almost a century to the works of Landau [1] and Pekar [2]. Formalised by Fröhlich [3], the polaron model is one of the most studied problems of quantum many-body theory [4] with real-world relevance for electronic material properties [5]. Calculations of the ground state energy, the dispersion relation and also the structure of the excitation continuum have been achieved at great accuracy and become an important test bed for analytical [6] and numerical [7, 8, 9, 10] techniques.
The existence of bound excited states below the excitation continuum in the strong coupling limit of the three-dimensional model was first predicted in the 1960s [11, 12], and rigorously proven very recently [13]. Previous numerical studies have failed to produce evidence for bound excited states in the three-dimensional case [8]. Studies of one-dimensional Fröhlich polaron models found excitations of the electron as resonances at high energies within the phonon continuum [14, 15]. In contrast, for the Holstein model of a small polaron, the existence of a bound excited state below the continuum is well known [16, 17, 18].
In this work we calculate the energies of the bound excited states of the phonon dressing cloud for the Fröhlich model in one dimension in the strong coupling limit, and provide numerical evidence for them in the weak to intermediate coupling regime applying Full Configuration Interaction Quantum Monte Carlo (FCIQMC) [19, 20, 21]. The first bound excited state appears at $\alpha\approx 1.73$, which is in a crossover regime between weak and intermediate coupling strength with a spectral weight of $\approx 0.2$, suggesting that these states may be accessible in experiments.
@@ I Introduction
II The model
The Fröhlich model [3] describes an electron in a polarizable medium. Medium polarization is predominantly provided by longitudinal optical phonons [34, 4], and the one-dimensional model applies to the highly anisotropic situation of co-linear motion of electron and phonon motion [35]. While the Coulomb interactions contributing to the coupling constant diverge for strictly one-dimensional systems [36, 37], their contributions are regularized for quasi-one-dimensional quantum wires and nanotubes [38, 39, 40]. We parameterize the one-dimensional model with the dimensionless coupling constant $\alpha>0$ in the form [35, 37]
@@ II The model
[…]
where $\hat{p}$ and $\hat{r}$ are the electron momentum and position, $\hat{a}^{\dagger}_{k}$ and $\hat{a}_{k}$ are the phonon creation and annihilation operators, $\omega_{\mathrm{LO}}$ is the frequency of longitudinal optical phonons (taken to be constant), and $m$ is the band mass of the electron. $L$ is the length of the system with periodic boundaries (eventually taken to be infinite), and $l_{0}=\sqrt{{\hbar}/{2m\omega_{\mathrm{LO}}}}$ is a natural unit of length.
@@
III Excitation energies in the strong coupling limit
In the limit of large $\alpha$, the discrete excited eigenvalues of $\hat{H}$ in Eq. (II) at total momentum $P=0$ can be obtained analytically, and are given by
@@ III Excitation energies in the strong coupling limit
Since $\omega_{j}<1$ for all $j$, we thus expect that the number of discrete eigenvalues below the continuum increases without bounds as $\alpha\to\infty$, and this can actually be proved rigorously [13].
As is well-known [44, 45, 46], the first term in Eq. (2), $\alpha^{2}/3$, represents the semiclassical energy obtained by minimizing the corresponding Pekar functional with respect to the electron wave function $\psi(x)$ and the classical polarization field $\varphi(x)$. The Pekar functional (given in dimensionless units integrating over $x=r/l_{0}$)
@@
is obtained as an expectation value of the Hamiltonian of Eq. (II) with respect to a product of the electron wave function $\psi(x)$ and a coherent state for the phonon field with amplitude $\varphi(x)$. The physical picture is that, at large coupling, the polarization field effectively becomes classical, and the electron occupies the ground state in the corresponding effective potential. The Pekar functional has the explicit minimizer $\psi^{\rm P}(x)=\sqrt{\alpha/2}\,\mathrm{sech}(\alpha x)$ for the electron wave function and $\varphi^{\rm P}(x)=\sqrt{2\alpha}\psi^{\rm P}(x)^{2}$ for the polarization field. The other terms in Eq. (2) involving the ($\alpha$ independent) frequencies $\omega_{j}$ quantify the quantum fluctuations of the field around its classical value [47, 48, 49, 50, 51, 13]. Letting the electron take the instantaneous ground state in a varying field, they are obtained by expanding $\mathcal{F}^{\rm P}(\varphi)=\min_{\psi}\mathcal{E}^{\rm P}[\psi,\varphi^{\rm P}+{\varphi}]$ to second order in ${\varphi}(x)$, with the result
@@
$\mathcal{F}^{\rm P}(\varphi)\approx-\frac{\alpha^{2}}{3}+\|\text{Im}({\varphi})\|^{2}+\langle\text{Re}({\varphi}),\hat{\omega}^{2}\text{Re}({\varphi})\rangle.$ (4)
$\mathcal{F}^{\rm P}(\varphi)$ is equal to the sum of $\int|\varphi|^{2}$ and the ground state energy of the effective Hamiltonian $\partial^{2}-2\sqrt{2\alpha}\,\mathrm{Re}(\varphi)$, where $\partial=d/dx$. Accordingly, simple second order perturbation theory allows us to express the Hessian operator as
$\hat{\omega}^{2}=1-8\alpha\psi^{\rm P}\frac{Q}{-\partial^{2}-2\sqrt{2\alpha}\varphi^{\rm P}+\alpha^{2}}\psi^{\rm P},$ (5)
where $\psi^{\rm P}(x)$ is understood as a multiplication operator, and $Q$ projects orthogonal to $\psi^{\rm P}$. Note that $\psi^{\rm P}$ is the zero-energy ground state of $\partial^{2}-2\sqrt{2\alpha}\varphi^{\rm P}+\alpha^{2}=A^{\dagger}A$ with $A=\partial-(\psi^{\rm P})^{\prime}/\psi^{\rm P}$. We can write $(A^{\dagger}A)^{-1}Q=A^{\dagger}(AA^{\dagger})^{-2}A$. Since $AA^{\dagger}=-\partial^{2}+\alpha^{2}$, and $A\psi^{\rm P}=\psi^{\rm P}\partial$, the Hessian operator further simplifies to
$\hat{\omega}^{2}=1+8\alpha\partial\psi^{\rm P}(-\partial^{2}+\alpha^{2})^{-2}\psi^{\rm P}\partial.$ (6)
@@
By scaling, the eigenvalues of $\hat{\omega}$ are independent of $\alpha$. We denote them by $0=\omega_{0}<\omega_{1}\leq\dots<1$. There is a single zero mode, $\omega_{0}=0$, resulting from translation invariance. The eigenvalue equation $\hat{\omega}^{2}f=\omega^{2}f$ can be rewritten in terms of $\lambda\equiv 1-\omega^{2}$ as
$\left(\partial^{2}-4\frac{(\psi^{\rm P})^{\prime}}{\psi^{\rm P}}\partial+4\alpha^{2}+\frac{8\alpha}{\lambda}(\psi^{\rm P})^{2}\right)f^{\prime}=0.$ (7)
This follows from the identity $\psi^{\rm P}(-\partial^{2}+\alpha^{2})^{2}(\psi^{\rm P})^{-1}=(\partial^{2}-4\frac{(\psi^{\rm P})^{\prime}}{\psi^{\rm P}}\partial+4\alpha^{2})\partial^{2}$. Changing variables from $x$ to $t=-\alpha^{-1}\psi^{\rm P}(x)^{\prime}/\psi^{\rm P}(x)=\tanh(\alpha x)$ and making the ansatz $f^{\prime}(x)=\ell(t)(1-t^{2})$ turns Eq. (7) into the standard Legendre equation
$(1-t^{2})\ell^{\prime\prime}-2t\ell^{\prime}+n(n+1)\ell=0,$ (8)
with $n(n+1)=2+4\lambda^{-1}$. The general solution is thus a linear combination of $L^{P}_{n}$ and $L^{Q}_{n}$, the Legendre functions. The condition that the resulting function $f$ be square-integrable introduces the constraints $\int_{-1}^{1}\ell(t)dt=0$ and $\int_{-1}^{1}\ell(t)\ln\frac{1+t}{1-t}dt=0$ or, equivalently, $\int_{-1}^{1}\ell(t)L_{0}^{P}(t)dt=0=\int_{-1}^{1}\ell(t)L_{0}^{Q}(t)dt$ (due to the fact that $\int_{-\infty}^{\infty}f^{\prime}(x)dx=0$ and also $\int_{-\infty}^{\infty}f(x)dx=0$ since $f$ itself is a derivative).
The even solutions of Eq. (8), satisfying the orthogonality constraints, are simply the Legendre polynomials $L^{P}_{n}$ for $n=2,4,6,\dots$. This leads to (with $j\equiv n-2$)
$\omega_{j}^{2}=1-\frac{4}{(j+4)(j+1)}\ ,\quad j=0,2,4,\dots$ (9)
In particular, $\omega_{0}=0$ and $\omega_{2}=\sqrt{7/9}$. The odd solutions are slightly more complicated. They are linear combinations of $L^{P}_{n}$ and $L^{Q}_{n}$ for non-integer $n$. Using the well-known integrals over products of Legendre functions in [52], the orthogonality conditions transform to the equation
$\psi_{\Gamma}(n+1)+\gamma=\frac{\pi}{2}\tan\frac{n\pi}{2}$ (10)
where $\psi_{\Gamma}=\Gamma^{\prime}/\Gamma$ denotes the digamma function, and $\gamma\approx 0.577$ is Euler’s constant. There are infinitely many solutions, one in every interval $(2j,2j+1)$ for $j\geq 1$, the smallest being $n_{1}\approx 2.523$. Denoting the solution in the interval $(2j,2j+1)$ by $n_{2j-1}$, we thus conclude that
$\omega_{j}^{2}=1-\frac{4}{(n_{j}+2)(n_{j}-1)}\ ,\quad j=1,3,5,\dots$ (11)
The lowest non-vanishing eigenvalues are given by $\omega_{1}\approx 0.647$ and $\omega_{2}=\sqrt{7/9}\approx 0.882$. The corresponding eigenstates alternate in parity, with the first excited state being even.
@@
From this expression we have to subtract the zero-point energy $1/2$ for every mode, since in the absence of the electron the field energy in Eq. (II) is normalized to have zero vacuum energy [i.e., $\sum_{k}\hat{a}^{\dagger}_{k}a_{k}=\frac{1}{2}\sum_{j\geq 0}(\hat{p}_{j}^{2}+\hat{q}_{j}^{2}-1)$]. This leads to a contribution to the ground state energy of the form $\tfrac{1}{2}\sum_{i\geq 0}(\omega_{i}-1)$, hence justifying Eq. (2) for $j=0$. Numerically, one finds $\tfrac{1}{2}\sum_{i\geq 0}(\omega_{i}-1)\approx-0.955$. For $j\geq 1$, the excitation energies correspond to excitations of different oscillators in Eq. (12). Since $\omega_{j}>\tfrac{1}{2}$ for $j\geq 1$, only single excitations can appear below the continuum. The zero mode $j=0$ needs to be discarded for the calculation of the excitation spectrum of Eq. (II) due to the restriction to zero total momentum.
IV Lee-Low-Pines transformation
Using a unitary transformation due to Lee, Low, and Pines amounting to a frame transformation [53], the electron coordinate dependence of the Fröhlich polaron Hamiltonian of Eq. (II) can be eliminated and the Hamiltonian rewritten in terms of the conserved total momentum $P$:
@@ IV Lee-Low-Pines transformation
$\quad-\hbar\omega_{\mathrm{LO}}\sqrt{\frac{2\alpha}{L/l_{0}}}\sum_{k}\left(\hat{a}^{\dagger}_{k}+\hat{a}_{k}\right).$ (13)
This equivalent Hamiltonian acts only on a Fock space of phonons for a given value of the total momentum $P$ and is more convenient for numerical calculations. We take the system size $L$ finite, which discretizes the phonon modes with the values of $k$ restricted to multiples of $2\pi/L$. A finite number of modes $M$ used in numerical calculations results in a momentum cutoff of $k_{c}=\pi(M-1)/L$. Using a larger number of modes increases the computational cost. Even so, the phonon Hilbert space dimension is still formally (and practically) infinite, as an arbitrary number of phonons can reside in each mode. Technically, the number of phonons in each mode is limited to be <256<256 by the 8-bit number type used for representing the Fock basis. We perform calculations with up to $M=73$ and Hilbert space dimension $256^{M}\approx 6\times 10^{175}$. In the following, we will restrict the analysis to $P=0$.
V FCIQMC
The algorithm samples the ground state of $\hat{H}_{\mathrm{LLP}}$ by repeatedly applying the operator $1-\text{d}\tau(\hat{H}_{\mathrm{LLP}}-S)$ to a coefficient vector. The shift parameter $S$ is updated to keep the norm (the number of ‘walkers’) approximately constant [23], and therefore provides an estimator of the eigenstate energy after equilibration. The operation is performed exactly for coefficients with a modulus above a given threshold and stochastically for the rest. Combined with stochastic vector compression [54], this implements a semistochastic algorithm, where part of the Hamiltonian matrix is treated deterministically while the rest is sampled stochastically [20, 27]. Crucially, contributions of opposite sign to a coefficient can annihilate and thus reduce sign noise. At sufficiently large walker number, the sign of the dominant coefficients becomes stable by a spontaneous symmetry-breaking process [30, 55] and the sign problem is mitigated. The residual noise leads to stochastic errors and also causes a small population control bias [56, 57, 24], which we control by increasing the walker number until undetectable levels of the bias are seen that are smaller than the stochastic error bars.
If both the box size $L$ and the momentum cutoff $k_{c}$ are large enough, the results obtained using the discretized model approach the continuum limit. We find that the energies converge exponentially in $L$, but only algebraically in $k_{c}$ (see Appendix A).
VI Ground state energy
In the weak coupling limit, the ground state energy has been found analytically using perturbation theory [58, 37] and the path integral formalism [35]. For $P=0$, the ground state energy is
@@ VI Ground state energy
Figure 1 shows the transition from the weak coupling regime, through an intermediate coupling regime around $\alpha\approx 2$ where both expansions approximately overlap, into the strong coupling regime as the coupling strength $\alpha$ increases. While there is no phase transition separating the weak and strong coupling regimes, different physical pictures apply [59]. Our FCIQMC approach allows us to probe both regimes and the transition region comfortably. At $\alpha=2$ we estimate the discretization error to be around 1% and about an order of magnitude larger than the stochastic error (see Appendix A).
VII Excited state energies
In order to calculate excited states numerically with FCIQMC, the computation is initialized with multiple starting vectors, which are then orthogonalized at every step (or a specified number of steps) of the computation using the Gram-Schmidt procedure. This ensures that the $n^{\text{th}}$ vector converges to the $n^{\text{th}}$ eigenvector, and the $n^{\text{th}}$ shift acts as an estimator for the $n^{\text{th}}$-lowest eigenvalue [21].
Even though the ground state calculation with the Hamiltonian (IV) is free of the Monte Carlo sign problem, excited state calculations introduce the sign problem [30, 31] due to the necessary existence of nodes in the excited state wave functions [21]. The sign problem manifests itself at low walker number in fluctuating signs of all state vector coefficients and an average value of the shift $\langle S\rangle$ that is lower than the correct energy [19, 33]. Increasing the walker number above a critical value, the shift average stabilizes and becomes independent of the walker number, indicating that the sign problem is overcome. For the studied range of $\alpha$ values we find that $10^{7}$ walkers is sufficient, indicating that the sign problem is relatively weak in this case [33]. This is remarkable as the unconstrained phonon occupation makes the Hilbert space of the Fröhlich Hamiltonian (IV) infinite. For details, see Appendix C.
Figure 2: Energies of excited states relative to the ground state for the one-dimensional Fröhlich model with zero total momentum. The grey shaded region shows the phonon continuum starting $1\hbar\omega_{\mathrm{LO}}$ above the ground state. The dashed horizontal line indicates the strong-coupling prediction for the first bound excited state from Eq. (11). Symbols and error bars show finite system numerical results calculated using FCIQMC for two different box sizes ($L=6l_{0}$, green, and $L=10l_{0}$, purple) and classified by parity (circles for even and crosses for odd parity). The states inside the phonon continuum are discretized and strongly affected by the finite box size. A discretization-insensitive even state with energy below the phonon continuum for $\alpha\gtrapprox 1.73$ provides numerical evidence for a bound excited state of the phonon dressing cloud. The inset shows a detail near the emergence of the bound state and includes additional data for box size $L=8l_{0}$.